OpenStudy (anonymous):

find all polar coordinates of point p where p = (9, -pi/5)

OpenStudy (anonymous):

p is a polar coordinate right?

OpenStudy (anonymous):

@tholn

OpenStudy (anonymous):

i think so @genius12

OpenStudy (anonymous):

\[(9,-\frac{\pi}{5}+2k\pi)\]might work

OpenStudy (anonymous):

they are asking for two @satellite73

OpenStudy (anonymous):

it says "all"

OpenStudy (anonymous):

R= 9.021906 t(in degrees) =3.9935205

OpenStudy (anonymous):

these are the answers (9, + 2nπ) or (-9,pi/5 + (2n + 1)π) (9, + 2nπ) or (9, pi/5 + (2n + 1)π) @satellite73

OpenStudy (anonymous):

or the ones that make most sense